OpenStudy (anonymous):

What is the solution set for this linear-quadratic system of equations? y = x^2 − x − 12 y − x − 3 = 0

OpenStudy (anonymous):

@welshfella @acxbox22

OpenStudy (anonymous):

{(-3, 0), (0, 3)} {(-3, 0), (4, 0)} {(-3, 0), (5, 8)} {(4, 0), (0, 3)}

OpenStudy (acxbox22):

first solve for y on the bottome equation y-x-3=0 y=x+3 now we can plug in (x+3) for y on the top equation

OpenStudy (acxbox22):

x+3=x^2-x-12 subtracting x+3 on the left side we get 0=x^2-2x-15 now factor that and solve for x

OpenStudy (anonymous):

So its c.

OpenStudy (acxbox22):

0=(x-5)(x+3) x=5,-3 now plug those two values into wither equation to get their respective y values